# Paper Quality Assessment based on Individual Wisdom Metrics from Open Peer Review

### v1

@@ 2.4 How things could be: an open framework for peer review
Based on the findings above, we propose a new open framework for scientific peer review (Figure 4) and assess its soundness in a computational model. In this framework, all papers are immediately published (Eisen et al.,, 2020). Post-publication, users on the platform self-select themselves to review the paper. In this framework, users may give quality ratings not only to papers but also to others’ reviews (Walther and van den Bosch,, 2012). The quality of the given publication is estimated from these reviews and the ratings of the reviews. Authors of new submissions, reviewers, and raters of reviewers all come from the same user pool.
Generative model
@@
We implement a generative model of this process (Figure 4A): The model consists of a set of papers and a set of reviewers. Each generated paper $i$ has a hidden ground truth quality $q_{i}\in(0,1)$. Each reviewer has a ground truth reviewer quality $p_{i}\in(0,1)$. Whenever reviewer $i$ writes a review for paper $j$, they assign it a score $\sim N(q_{j},\alpha/p_{i})$: in other words, the score is based on the intrinsic paper quality, with a standard deviation that is inversely proportional to reviewer quality. Whenever user $i$ rates a review of another user $j$, they assign a score to them which is $\sim N(p_{j},\alpha/p_{i})$. We set $\alpha=0.18$ to match the reviewer correlation levels to those we found in CCN2023 data. We set the standard deviations of the review scores and paper scores to be equal, based on the finding from a NeurIPS 2022 randomized control trial (Goldberg et al.,, 2024) that disagreement between reviewers assessing review quality are comparable to the disagreement rates of paper reviewers. Varying $\alpha$ does not change the results qualitatively. In addition, we consider the existence of consistently low-quality, unreliable reviewers or “bots”. These unreliable users are modeled as assigning paper review scores and ratings of other reviews uniformly at random in$[0,1]$, regardless of paper or review quality.
@@
All assigned scores must lie in the interval $[0,1]$, and user-assigned scores are the only data available for the system to derive its quality estimators. As a worst-case conservative scenario, we consider a batch/offnline process, in which users have provided reviews of papers and other users, and we must estimate all quantities (paper quality and reveiwer quality) at the same time: scoring is based on all available papers and reviews available at the instantaneous time. (A history-dependent process, which we discuss below, may enable further accuracy gains.) We further assume, as a conservative poor-case scenario and consistent with the CCN 2023 dataset, that we are in a regime where assessing quality is hardest: the low-review regime in which each user publishes one paper a year, reviews $\leq$ 3 papers and rates $\leq$ 10 other users, meaning that each paper receives 3 reviews on average.
Scoring
@@
Scoring proceeds as follows: a rating is calculated for each reviewer by averaging all scores that this reviewer received for their reviews from other users. We then bin reviewers by their reviewer rating, and estimate how closely each bin predicts the CAS for any submission (Figure 4B). We also estimate the mean squared deviation (MSD) of each bin’s scores around the CAS, on the subset of papers on the platform for which the CAS is known with high certainty (in this example we use the top 20% reviewed papers in terms of the number of reviews received, when only including reviews by reviewers which have a rating in the top 20% percentile).
@@
Computing the MSD of reviewers allows for the generation of an estimation certainty score for each paper, based on the equation (derived in Appendix 1)
$1/\sigma^{2}_{\text{total}}=\sum 1/\sigma_{i}^{2}.$ (3)
This weighting procedure, combined with only committing to publish the scores that pass a certain threshold of certainty (here – $\sigma_{\text{total}}<0.15$) results in an estimator of paper quality (Figure 4C; solid black line) that tracks the (hidden) ground truth quality of papers on the platform more closely than the simple mean estimator (Figure 4C; solid gray line).
Unreliable reviewers
@@
Unreliable reviewers, defined for our purposes as users who assign random reviews that are uncorrelated with user and paper quality, may consist of different populations: they could be scientists who do not invest time or effort in the review process, or they could be adversarial human or even non-human (”bot”) actors.
@@
The standard way to reduce the influence of non-human bots is to require all users who deposit papers or review papers to be verified scientists (Walther and van den Bosch,, 2012). Such a step would be essential because the number of non-human bots can overwhelm the number of human users by orders of magnitude. After this step, in the low-review regime with the possibility of unreliable human users, imposing a constraint on minimum certainty before publishing a score results in substantial robustness to contamination with noise and unreliable reviewers, as done in Figure 4C.
@@
In addition, given that there are two qualitatively distinct types of reviewers in the system (those who issue scores with some correlation to paper quality and those whose outputs are uncorrelated), we consider entirely thresholding out (removing from consideration) the bottom percentiles of reviewers instead of weighting their reviews by the Bayesian MSD review quality factor, reasoning that this bin will contain mostly the uncorrelated reviewers. When we compute the correlation of the resulting paper quality scores with ground-truth paper quality, we find that, surprisingly, thresholding reviewers by their reviewer quality does not yield a prediction improvement over the full Bayesian weighting (Figure 4C, solid red line). As a sanity check, if we directly use a reviewer quality ”oracle” estimator, based on directly estimating the correlation of paper scores assigned by the reviewer to ground truth paper scores, and threshold reviewers out based on this oracular reviewer quality estimate, our reasoning holds and thresholding significantly improves paper quality estimation (Figure 4C, dashed black line).
@@
The cause of the discrepancy between the oracular review quality thresholding and the normally estimated review quality thresholding result becomes clear when we examine the distributions of estimated reviewer quality, Figure 4D: through the unreliable reviewers have a ground-truth quality of 0, as they become a larger fraction of the reviewing population, their reviews lead other unreliable reviewers to have a broad distribution of reviewer quality scores, meaning they are no longer confined to the lower quality bins. Thresholding out the lowest reviewer quality bins does not target only or the majority of the actual low-quality reviewers. This result suggests that it might be possible to further improve the fidelity of review with a “warm start” or online reviewing process, in which the platform begins with high-quality reviewers. New reviewers can join, but they must work their way up (in the sense of obtaining above-threshold review quality scores from estabilished users) for their reviews to influence paper quality scores. Future work will examine a model that includes such a warm-start process.
Reviewer Quality Incentives
In addition to improving estimates of paper quality on the platform, we expect the quantification of reviewer accuracies to incentivize reviewers to write high-quality reviews, because it would place better reviewers in a higher reviewer quality bin, and weigh their decisions about future papers more heavily (Figure 5A). The inverse of the reviewer MSD is also a reviewer quality score that can be used to recognize not only the work of the reviewer but also the quality of their work. The reviewer quality score could be used to provide reviewer credit in different ways, including for promotion or tenure decisions, for public display via a quality certificate issued to the reviewer, and so on. This recognition and accountability would provide a second incentive for writing higher-quality reviews (Pascual-Ezama et al.,, 2015).
Paper Coverage Incentives
The estimation certainty measure (Eq. 3) can be used to incentivize reviewers to provide better coverage for papers whose quality certainty is low. In a simple experiment, we consider that reviewers select papers to review based on the Chinese restaurant process prior (CRP), a stochastic process where the probability of a paper being selected is proportional to how many reviews the paper already has, in other words its popularity (a typical rich-get-richer process; Figure 5B, left). Next, we define a reward signal, which we set to be equal to $\Delta\sigma^{2}_{\text{total}}$ the amount of reduction in the uncertainty in the paper’s score that can be provided by this reviewer. Reviewers are rewarded for reviewing papers for which they can provide a strong boost in score estimation certainty. We assume that through offering platform incentives and rewards, the Chinese restaurant process probability of a reviewer choosing to review a certain paper can be modulated by this reward score. We find that such a process can avoid the typical snowballing dynamics of popularity that are evident in many social media systems (Narayanan,, 2023), and lead to a broader coverage of reviewed papers across the whole simulated platform (Figure 5B, right).
@@
Figure 5: Embedding incentives for consistently writing high-quality reviews and ensuring broad coverage within the open platform framework. (A) Left: Bayes weighting allows reviewers with higher ratings have more influence on the final paper score estimation, which provides one type of incentive for providing quality reviews. Right: Users can publicly advertise and use for promotion and tenure their reviewer ratings, another incentive for high-quality review. (B) Left: Distribution of reviews per paper, assuming papers accrue reviews according to a Chinese restaurant process (CRP). Right: Rewarding reviewers based on their contribution to reducing the uncertainty in the reviewed paper’s score estimate can provide , typical snowballing dynamics of popularity, evident in any form of social media, may be softened or avoided.
@@ 3 Discussion
We evaluated two real-world peer-review datasets to determine how well reviewer scores correlate with each other, and found very low levels of agreement between reviewers. In one of these datasets, we made an estimate of paper quality for a subset of papers that had a large number of reviews based on the community average score, and used this as a metric for assessing both authorship quality and reviewer quality for each individual reviewer. We found that the authorship quality of an individual is not predictive of their reviewer quality.
Given these quantified challenges in extracting meaningful paper quality scores from small numbers of noisy reviewers, we proposed and assessed a Bayesian weighting of reviewer scores based on empirically estimating reviewer quality. In the low number of reviews per paper regime, the proposed measure significantly outperforms standard averaging methods. We mitigated the problem of estimating reviewer quality given the small numbers of reviews written by each individual with a method for binned estimation of reviewer quality.
Finally, we showed that generating reviewer scores can incentivize reviewers to produce high-quality reviews in two ways: a desire for impact (a reviewer’s assessment of a paper is weighted more heavily if they have a higher reviewer score), and a desire for recognition (a reviewer’s work can be visibly recognized in the form of a reviewer score that can be used for promotion and tenure decisions and as a badge of honor). On top of such an incentive structure, we explored the possibility of incentivizing reviewers to cover a larger set of papers, avoiding the rich-get-richer phenomenon in which a few papers get heavily reviewed while most are neglected.
@@
Open peer review frameworks for scientific publishing, where reviews can be made public and evaluated, have been widely proposed over the past decade (Eisen et al.,, 2020; Ginsparg,, 1997; Eisen,, 2016; Kravitz and Baker,, 2011; Kriegeskorte et al.,, 2012; Nosek and Bar-Anan,, 2012; Stern and O’Shea,, 2019; da Silva,, 2013; LeCun,, 2013; Walther and van den Bosch,, 2012). The general idea of users generating commentary and then allowing users to rate the commentary of others has been implemented to considerable success in platforms like StackOverflow and Reddit, with its drawbacks investigated (Davis and Graham,, 2021; Mazloomzadeh et al.,, 2024; Melnikov et al.,, 2018; Movshovitz-Attias et al.,, 2013; Wang et al.,, 2021). Specifically in the field of scientific peer review, Goldberg et al., (2024) presented a randomized control trial, carried out at NeurIPS 2022, which showcases the measures of inter-evaluator disagreement, miscalibration, subjectivity, and biases in peer reviews of peer reviews. Rating user comments is considered a low-bandwidth and simpler exercise than generating commentary. We find that it provides a useful signal for open peer review that simultaneously enables positive incentives as well as more accurate community belief estimation. Recommendation systems in general and specifically for disseminating scientific research are likewise being investigated (Isinkaye et al.,, 2015; Bai et al.,, 2020; Achakulvisut et al.,, 2016; Ko et al.,, 2022; Pinedo et al.,, 2024; Putra Utama et al.,, 2023; Shani and Gunawardana,, 2011). In general settings, crowd-based metrics have been proposed to extract the wisdom of the crowd from the noisy and sometimes on average incorrect signal (Prelec et al.,, 2017; Kameda et al.,, 2022; Lee,, 2024; Palley and Soll,, 2019).
Outlook
@@
The volume of scientific publication has been increasing exponentially, and the difficulty of securing accurate high-quality paper assessments, in a timely way, has grown in tandem (Shah et al.,, 2017; Kuznetsov et al.,, 2024). The scientific community has shied away from bottom-up peer review to mitigate the potential for low-quality assessments. We conceive of the process as a dynamical system, in which the right mixture of estimation and incenitives can produce self-organized high-quality reviewing ecosystems. Here, we have taken small steps to explore possible metrics and incentives. Implementing and testing these ideas in open publishing platforms should help to provide real-world tests and improvements of such methods, possibly leading to a new process for scientific publishing.

### v2

@@ 2.4 Open peer review with user-generated ratings of both papers and reviews
In the hypothesized framework, all papers are immediately published (Eisen et al.,, 2020). Post-publication, users on the platform self-select themselves to review the paper. Users may give quality ratings not only to papers but also to others’ reviews (Walther and van den Bosch,, 2012). The quality of the given publication is estimated from these reviews and the ratings of the reviews. Authors of new submissions, reviewers, and raters of reviewers all come from the same user pool. In the subsections below, we assess the reliability of this process.
Figure 4: Reviewing reviews: Benefits of reviewer quality metrics in a richer ecosystem. (A) An open platform where content is published immediately, reviewers are self-selected, and the community assesses (rates) review quality. (B) Estimation of reviewer accuracy: binning reviewers by the ratings their reviews received enables estimation of that group’s accuracy in paper evaluation (mitigating the problem of estimating an accuracy score based on individual reviewer scores, given the small number of reviews supplied by each reviewer). (C) Bayes weighting of review scores based on reviewer accuracy as estimated from (B) leads to a better estimate of the ground truth paper quality scores (left; solid black line) than simple averaging of all review scores (left; gray line). An alternative heuristic of simply thresholding out all reviewers with a sufficiently low rating percentile (here, below 80 percentile, red line) results in a similar correlation to Bayes weighting but much lower coverage of papers on the platform (right). (D) Distributions of estimated reviewer quality scores for bots and real users in the simulated platform. In high-noise regime with most of the users being bots (up to 80%), reliability of the reviewer quality estimation breaks down, suggesting that alternative ways of excluding bots might be more effective.
Methods: Generative model
@@
We implement a generative model of this process (Figure 4A): The model consists of a set of papers and a set of reviewers. Each generated paper $i$ has a hidden ground truth quality $q_{i}\in(0,1)$. Each reviewer has a ground truth reviewer quality $p_{i}\in(0,1)$. Whenever reviewer $i$ writes a review for paper $j$, they assign to it a score $\sim N(q_{j},\alpha/p_{i})$: in other words, the score is based on the intrinsic paper quality, with a standard deviation that is inversely proportional to reviewer quality. Whenever user $i$ rates a review of another user $j$, they assign a score to them which is $\sim N(p_{j},\alpha/p_{i})$. We set $\alpha=0.18$ to match the reviewer correlation levels to those we found in CCN2023 data. We set the standard deviations of the review scores and paper scores to be equal, based on the finding from a NeurIPS 2022 randomized control trial (Goldberg et al.,, 2024) that disagreement between reviewers assessing review quality are comparable to the disagreement rates of paper reviewers. In our experiments, arying $\alpha$ did not change the results qualitatively. In addition, we consider the existence of consistently low-quality, unreliable reviewers or “bots”. These unreliable users are modeled as assigning paper review scores and ratings of other reviews uniformly at random in $[0,1]$, regardless of paper or review quality.
@@
All assigned scores must lie in the interval $[0,1]$, and user-assigned scores are the only data used by the system to derive its quality estimators. As a conservative scenario, we first consider a batch/offline process, in which users have provided reviews of papers and other users, and we must estimate all quantities (paper quality and reviewer quality) at the same time: scoring is based on all available papers and reviews available at the end of the review process (a history-dependent process, which we discuss below, may enable further accuracy gains). We further assume, as a conservative poor-case scenario and consistent with the CCN 2023 dataset, that we are in a regime where assessing quality is hardest: the low-review regime in which each user publishes one paper a year, reviews $\leq$ 3 papers and rates $\leq$ 10 other users, meaning that each paper receives 3 reviews on average.
Methods: Scoring
@@
Scoring proceeds as follows: a rating is calculated for each reviewer by averaging all scores that this reviewer received for their reviews from other users. We then bin reviewers by their reviewer rating, and estimate how closely each bin predicts the CAS for any submission (Figure 4B). We also estimate the mean squared deviation (MSD) of each bin’s scores around the CAS, on the subset of papers on the platform for which the CAS is known with high certainty (in this example we use the papers with most number of reviews received – top 20% – when only including reviews by reviewers who have a rating in the top 20% percentile).
@@
Computing the MSD of reviewers allows for the generation of an estimation certainty score for each paper, based on the equation $1/\sigma^{2}_{\text{total}}=\sum 1/\sigma_{i}^{2}$ (Appendix 1; Eq. 2). We combine this weighting procedure with committing to publish only those scores that pass a certain threshold of certainty (here – $\sigma_{\text{total}}<0.15$).
Direct user ratings of reviews improves paper scoring
The result of the procedure above is an estimator of paper quality (Figure 4C; solid black line) that tracks the (hidden) ground truth quality of papers on the platform more closely than the simple mean estimator (Figure 4C; solid gray line). It is possible to create an estimator of paper quality through a procedure that interpolates between a simple mean and the Bayesian estimator, by considering and averaging together (without further weighting) only the reviews submitted by the top 20% of reviewers for each paper. The performance of this estimator on a given paper matches that of the Bayesian estimator (Figure 4C; solid red line), but has the downside that its coverage of papers is sharply reduced, with assessments available for a much smaller number of papers (only 45% compared to 60% of papers with the Bayesian method; Figure 4C, right). Thus, our Bayesian estimator makes more efficient use of existing reviewers, including poorer reviewers, to arrive at good quality assessments of a larger fraction of papers.
We compared the performance of the Bayesian estimator above, which is based on direct user evaluation of review quality with Bayesian estimation performance, with the simpler reviewer quality estimation procedure (from Figure 2B and the CCN data analysis of Figure 3), which ignores the scores given by users to reviews and estimates each reviewer’s S.D. directly based on the deviations of their reviews from the CAS. The performance of this procedure depends heavily on the mean number of reviews each paper receives: in a typical situation of 3 reviews per paper, it performs even worse than simply averaging of all review scores (Figure 4E, left). When the number of reviews approaches that of CCN data (around 9), performance is slightly better than the simple mean (Figure 4E, middle), consistent with our finding in real CCN data (Figure 3). As the number of reviews grows to 20, the performance of the direct S.D. estimator matches that of our best proposed measure from the above subsection (Figure 4E, right). However, there are two serious drawbacks: this procedure should not naively be implemented in any online peer review platform, as it rewards reviewers who have small score deviations from CAS, encouraging reviewers to give papers the same score as all the other reviewers. This incentive can lead to a collapse scenario where each reviewer simply selects whatever rating was popular among other reviewers. And, even post-factum, this method only matches the direct user evaluation of reviewer-quality when the number of reviews is large.
Reducing unreliable reviewers
@@
Unreliable reviewers, defined in our model as users who assign random reviews that are uncorrelated with user and paper quality, may map in reality to different populations: they could be scientists who invest minimal time or effort in the review process, or they could be adversarial human or even non-human (”bot”) actors.
@@
One way to reduce the influence of non-human bots is to require all users who deposit papers or review papers to be verified scientists (Walther and van den Bosch,, 2012). Such a step is essential because the number of non-human bots can overwhelm the number of human users by orders of magnitude. After this step, in the low-review regime with the possibility of unreliable human users, imposing a constraint on minimum certainty about the resulting paper score before publishing a score results in substantial robustness to contamination with noise and unreliable reviewers, as done in Figure 4C. Note that the minimum certainty constraint is required to achieve the performance gains, even when using the oracle with perfect information about reviewer quality (Supplementary Figure S4).
@@
In addition, given that our generative model includes two qualitatively distinct types of reviewers (those who issue scores with some correlation to the underlying paper quality and those who issue pure noise), we consider entirely thresholding out (removing from consideration) the bottom percentiles of reviewers instead of weighting their reviews by the Bayesian MSD review quality factor, reasoning that this bin will contain mostly the uncorrelated reviewers.
Indeed, if we could directly use a reviewer quality ”oracle” estimator, which would have perfect information about the distribution of each reviewer’s deviations from CAS, including thresholding out bots, we could achieve a much higher paper quality estimation (Figure 4C, dashed black line). However, in the real-world case where ground truth paper scores are unknown, we find that, surprisingly, thresholding reviewers by their estimated reviewer quality (Figure 4C, solid red line) does not yield a prediction improvement over the full Bayesian weighting (black line).
@@
The cause of the discrepancy between the oracular review quality thresholding and the normally estimated review quality thresholding result becomes clear when we examine the distributions of estimated reviewer quality, Figure 4D: even though the unreliable reviewers have a ground-truth reviewer quality of 0, as they become a larger fraction of the reviewing population, their reviews lead other unreliable reviewers to achieve a broad distribution of reviewer quality scores, meaning they are no longer confined to the lower quality bins. Thresholding out the lowest reviewer quality bins does not target exclusively or even a majority of actual low-quality reviewers. This result suggests that it might be possible to further improve the fidelity of review with a “warm start” or online reviewing process, in which the platform begins with high-quality reviewers. As new reviewers join, they must work their way up (in the sense of obtaining above-threshold review quality scores from established users) for their reviews to influence paper quality scores and reviewer scores. We investigate this scenario in a subsection below.
Binary (up/down) scoring of review quality does not substantially degrade performance
In this section, we investigate the effects of binarizing the scores of others’ reviews on the platform (binary scores). We use the same model from Figure 4, which runs for 5 “years.” On average, each user writes 3 reviews and provides 10 scores per year (or 50 scores in the 5X scores condition). For this analysis, we fix the proportion of bots on the platform to be $50\%$. We find that there is a borderline but non-significant difference between performance of the Bayes-weighted measure between the baseline and binary scores condition ($P\approx 0.05$), Fig. 5, suggesting that the platform does not suffer from switching from real number scores in $[0,1]$ to simpler binary scores which might require less effort from the reviewers. Furthermore, we conjecture that the ease of assigning a binary score will incentivize users to provide more of those scores, improving the overall platform performance; and indeed, we notice a significant increase in the correlation between estimated and real paper quality for the Bayes-weighted metric when increasing the number of scores of others’ reviews by the factor of 5 (while keeping the number of reviews per user the same).
Figure 5: Binarized scoring of reviews does not reduce performance. Left: The baseline condition, where all scores and reviews are real numbers between 0 and 1. Center: The ”binary scores” condition, where the reviews of reviews (but not paper review scores themselves) are binarized using a threshold of 0.5, does not strongly reduce performance. Right: The binary scores condition, where the number of binary review of reviews scores are 5 times as numerous as on the left (under the assumption that it’s significantly easier to provide a binary label than a real score), leads to improved performance relative to fewer non-binary scores, ***p<<0.0001.
Figure 6: “Warm start” platform initialization can result in better performance. Left: Performance in the baseline condition, where the initial reviewers are randomly picked from the pool. Right: Warm start condition: starting from the same pool as in the left, initial reviewers are picked (without replacement) to be the best available reviewers. Subsequent reviewers are selected randomly from the remainder of the pool. Results are for simulations over 1 (A), 3 (B), or 5 (C) years, with 80% of the users designated as bots. The warm start advantage persists over years but gradually decreases.
“Warm Start”
Next, we investigate the effects of initializing the platform with the best available reviewers, a process we will call a warm start. We consider how this choice might change the long-term dynamics on the platform. We create simulations using the same model as Figure 4A, running them for multiple “years”. In each year, we simulate the onboarding of new users, content creation, and user exits. Specifically, the simulation begins with 500 users and 20 content items (papers) per user. Each year, 2000 new users join the platform, while 10% of existing users exit. Users do not generate new content after onboarding, but they do contribute reviews: on average, each user writes 3 reviews and provides 10 scores per year. Here we only quantify an unthresholded warm start (everyone gets a vote, but the platform is seeded with good reviewers), rather than the more aggressive policy of initially assigning zero weights to incoming reviewers until their reviewer quality reaches a threshold.
To assess the value of a warm start process, we compare two scenarios: one in which the initial 500 users are drawn randomly from the population, and one in which they are selected (without replacement) from the top-performing reviewers (Figure 6A-C). In both cases, the distribution of review qualities in the user pool is the same; the only difference is the ordering. Results show that warm-starting with high-quality reviewers leads to significantly higher correlations between the estimated and true paper quality, and the effect is largest for the Bayes-weighted measures. The effects are largest in early years and decay gradually over time, but with a persistence of several years.
2.5 Extrinsic reviewer incentives
There are many reasons why scholars participate in peer review: Altruistic contribution to one’s community and the chance to shape the content, claims, citations, and messaging of papers. These intrinsic incentives exist without the community granting overt recognition of the contributions made by reviewing, and without specific incentives delivered to review particular papers. Additionally, in the system outlined above, reviewer quality and thus their impact (weighting) on a paper’s assessment is determined based on their quality score, not on the number of reviews they have written (Figure 7A). Thus, there is an natural bias towards quality over quantity.
Here we briefly consider additional extrinsic incentives that could further improve reviewer participation, quality, or distribution over papers.
Reviewer quality publicization
The computed reviewer quality score can be used to explicitly recognize the quality of a reviewer’s work, in the form of a quality certificate. This certificate could be generated publicly for each reviewer without attaching their anonymous reviewer name, thus preserving review anonymity while providing public acknowledgment. Quality certification could be used in various ways, including for promotion or tenure decisions, for public display on websites, and so on. This recognition and accountability would provide additional extrinsic incentives, beyond the intrinsic ones, for writing higher-quality reviews (Pascual-Ezama et al.,, 2015).
Scarcity incentives
Some reviewers may be inclined to write vastly more paper reviews than others, diluting the impact of the rest. We propose explicitly capping the number of reviews than can be written by each user per year, with a possible bonus of 1-2 additional reviews allowed to reviewers depending on their quality score. This makes explicit that quality not quantity determines reviewer scores, and guards against crowd effects in the review process. In addition, as scarcity creates value, we believe that restricting the ability to review will make reviewing a more valued activity. We encourage future modeling work to take into account human psychology around scarcity to determine the quantitative effects of review caps on reviewer participation and quality.
Incentive to broaden review coverage across papers
A common dynamic in nature including human social interactions is the rich-get-richer process, one in which popular or large items become more popular or larger, leading to heavy tails in various distributions. These dynamics are evident in “viral” content and viral content creators on many social media systems (Narayanan,, 2023). In the context of bottom-up peer review, a concern is that a few papers might come to garner the lion’s share of reviewer attention while many are ignored. Though capping the number of reviews each individual can contribute partially mitigates this effect, it is still possible for a large fraction users to decide to review an already “viral” paper.
We consider an information-gain metric to provide further incentives for broad coverage of reviews. Our Bayesian paper evaluation approach process naturally generates an estimation certainty measure per paper (Eq. 2), which indicates which papers require more reviews. In a preliminary experiment, we assume that the user process of selecting which papers to review is based on the so-called Chinese restaurant process (CRP), in which the probability of selecting a paper is proportional to how many reviews the paper already has ($p\propto\#\text{reviews}$, a common dynamic in human group behavior; Figure 7B, left). With this process, which we treat as the baseline condition, a highly skewed popularity distribution emerges, in which a small minority of papers amass most reviews (Figure 7B, left). Next, we define a reward signal for each paper, equal to $\Delta\sigma^{2}_{\text{total}}$ – the amount of reduction in the uncertainty in the paper’s score that would be provided by this reviewer were they to review the paper. We assume that the reward modifies reviewer behavior by convolving the CRP with the reward signal, $p\propto\#\text{reviews}\cdot\Delta\sigma^{2}_{\text{total}}$. When reviewer behavior is modulated by the uncertainty-reduction reward in this way, the result is a substantially broader coverage of reviewed papers across the simulated platform (Figure 7B, right), avoiding the rich-get-richer phenomenon. This suggests that providing reviewers with additional incentives based on the uncertainty-reduction metric could be a powerful tool to obtain a broad reviewer coverage of papers.
@@ 2.5 Extrinsic reviewer incentives
Figure 7: Embedding incentives for writing high-quality reviews and ensuring broad coverage within the open platform framework. (A) Left: Bayes weighting awards high-quality reviewers with more influence on the final paper score estimation, providing one type of incentive for providing quality reviews. Right: In addition, users could publicly show and use for promotion and tenure their reviewer ratings, another incentive for high-quality review. (B) Left: Distribution of reviews per paper, assuming reviewers select papers according to a Chinese restaurant process (CRP). Right: Rewarding reviewers based reducing the uncertainty in the reviewed paper’s quality estimate can help flatten the distribution.
@@ 3 Discussion
This paper makes several contributions: we evaluated two real-world peer-review datasets to determine how well reviewer scores correlate with each other, and found very low levels of agreement between reviewers. Using one of the datasets, we showed that there is little correlation between authorship quality and reviewer quality (by computing the community averaged score for paper quality for a subset of papers with a large number of reviews and using it as a metric to assess both authorship quality and reviewer quality per reviewer), so that the authorship quality of an individual is not predictive of their reviewer quality.
Given these and other quantified challenges in extracting meaningful paper quality scores from small numbers of noisy reviewers, we proposed and assessed Bayesian weighting methods to estimate paper quality based on 1) an empirical estimate of reviewer quality from paper review data or 2) a direct additional evaluation of reviewer quality by other reviewers. In the low number of reviews per paper regime, we showed that Bayesian paper scoring with both proposed reviewer estimation methods significantly outperform standard averaging. In the first method, we mitigated the problem of estimating reviewer quality given the small numbers of reviews written by each individual with a method for binned (collective) estimation of reviewer quality. In addition, we showed that using direct and noisy but unbiased reviewer/reader feedback to determine reviewer quality (method 2) is better than indirect estimation of reviewer quality (method 1) paper quality estimation, even when user feedback has low bandwidth, in the form of binary feedback (simple up- or down-voting). Thus, we have shown through analysis of open review data and modeling that there are simple and effective ways to mitigate the problem of small-$N$ noisy reviews in estimating paper quality, even with a majority of completely unreliable but unbiased agents. An important direction for future work is to investigate whether combining both reviewer quality estimation methods (the empirical estimate based on deviation from community average scores and the direct peer evaluation of reviews) could yield further improvements in paper quality estimation.
Finally, we explored and quantified the potential of specific metrics as incentives to avoid rich-get-richer snowballing phenomena in paper coverage by reviewers. We showed that an information gain metric could, if implemented as an incentive that influenced reviewer choice, lead to much broader and even coverage of the set of available papers.
@@
Open peer review frameworks for scientific publishing, where reviews can be made public and evaluated, have been widely proposed over the past decade (Eisen et al.,, 2020; Ginsparg,, 1997; Eisen,, 2016; Kravitz and Baker,, 2011; Kriegeskorte et al.,, 2012; Nosek and Bar-Anan,, 2012; Stern and O’Shea,, 2019; da Silva,, 2013; LeCun,, 2013; Walther and van den Bosch,, 2012). The general idea of users generating commentary and then allowing users to rate the commentary of others has been implemented to considerable success in platforms like StackOverflow and Reddit, with its drawbacks investigated (Davis and Graham,, 2021; Mazloomzadeh et al.,, 2024; Melnikov et al.,, 2018; Movshovitz-Attias et al.,, 2013; Wang et al.,, 2021). Specifically in the field of scientific peer review, Goldberg et al., (2024) presented a randomized control trial, carried out at NeurIPS 2022, which showcases the measures of inter-evaluator disagreement, miscalibration, subjectivity, and biases in peer reviews of peer reviews. Rating user comments is considered a low-bandwidth and simpler exercise than generating commentary. We find that it provides a useful signal for open peer review that simultaneously enables positive incentives as well as more accurate community belief estimation. Recommendation systems in general and specifically for disseminating scientific research are likewise being investigated (Isinkaye et al.,, 2015; Bai et al.,, 2020; Achakulvisut et al.,, 2016; Ko et al.,, 2022; Pinedo et al.,, 2024; Putra Utama et al.,, 2023; Shani and Gunawardana,, 2011). In general settings, crowd-based metrics have been proposed to extract the wisdom of the crowd from the noisy and sometimes on average incorrect signal (Prelec et al.,, 2017; Kameda et al.,, 2022; Lee,, 2024; Palley and Soll,, 2019). We believe that the present work, based on concrete suggestions for improving peer assessment and evaluation through generative modeling and analysis of open review data, adds an actionable contribution for the future design and implementation of new systems.
Outlook: How things could be
The volume of scientific publication has been increasing exponentially, and the difficulty of securing accurate high-quality paper assessments, in a timely way, has grown in tandem (Shah et al.,, 2017; Kuznetsov et al.,, 2024).
Generating, using, and making available lifetime reviewer scores can incentivize scholars to write reviews and focus on quality, the core challenge for potential open review platforms. We believe this to be the case for two reasons related to innate scholarly drive: the first is a desire for impact (a reviewer’s assessment of a paper is weighted more heavily if they have a higher reviewer score), and the second is a desire for recognition (a reviewer’s work can is visibly quantified and acknowledged in the form of a score/certification that can be used for promotion and tenure decisions and as a badge of honor).
@@
The scientific community has shied away from bottom-up peer review to mitigate the potential for low-quality assessments. We conceive of the process as a dynamical system, in which the right mixture of estimation and incenitives can produce self-organized high-quality reviewing ecosystems. Here, we have taken small steps to explore possible metrics and incentives. Implementing and testing these ideas in open publishing platforms should help to provide real-world tests and improvements of such methods, possibly leading to a new process for scientific publishing.